# From Word Vectors to Multimodal Embeddings: Techniques, Applications, and Future Directions For Large Language Models

### v1

@@ IV-C2 Methods for Modeling Brain Activity
Several neuroimaging techniques allow researchers to study brain activity during language processing. Functional magnetic resonance imaging (fMRI) measures brain activity through blood flow changes, providing insight into the brain regions involved in various language tasks. [72] used fMRI to analyze the impact of contextual information on brain responses. Electroencephalography (EEG) captures electrical activity from scalp electrodes, offering high temporal resolution for tracking neural response timing to language stimuli. [73] applied EEG to examine brain responses to familiar entities. Electrocorticography (ECoG), which involves placing electrodes on the brain’s surface, provides even higher spatial resolution than EEG. [74] utilized ECoG to study the geometric similarity between brain embeddings and contextual language model embeddings. These techniques capture neural responses to linguistic stimuli, ranging from single words to complex narratives.
@@
Brain embeddings are continuous vector representations derived from neural activity in response to linguistic stimuli. These embeddings encapsulate patterns of brain responses across specific time windows and regions of interest (ROIs). For example, [74] extracted brain embeddings by averaging neural activity in the inferior frontal gyrus (IFG) during a listening task. This enables direct comparison with word embeddings derived from language models, facilitating a deeper understanding of how the brain processes language.
@@
Techniques such as geometric alignment and zero-shot mapping are used to compare brain embeddings with language model embeddings. Geometric alignment assesses the structural similarity between embedding spaces, while zero-shot mapping involves predicting brain activity patterns for new words based on their language model representations, or vice versa. [71] adapted a transfer learning approach to reveal a low-dimensional structure in language model representations that aligns with brain responses. [74] found that brain embeddings and DLM embeddings share similar geometry, enabling zero-shot predictions of brain activity for novel words, suggesting a potential alignment between human and artificial language representation.
@@ IV-C3 Relating Language Model Layers to Temporal Dynamics of Brain Activity
Research has explored the relationship between the layered architecture of deep language models (DLMs) and the temporal dynamics of brain activity. [75] found a correspondence between sequential transformations in DLM layers and the temporal sequence of neural activity in brain regions associated with language. This suggests that different DLM layers may reflect different stages of language processing in the brain. However, the study also identifies unique neural processes, such as recurrent information accumulation and responses to unpredictable words, which are not fully captured by current DLMs.
@@
While language models are typically trained on large general corpora, individual experiences shape human language understanding. [73] demonstrated that language models could represent subject-specific semantic knowledge of familiar people and places when trained on personalized data. This capability opens up possibilities for personalized language models that capture both general and individual semantic knowledge, with applications in personalized recommendations and user-specific content generation.
@@
Despite recent advancements, challenges remain in modeling brain activity with language model embeddings. Understanding the dimensionality of brain embedding spaces and how they relate to language model dimensions is essential. Another key challenge is modeling neural processes for unpredictable words, which current models do not fully capture. Additionally, linking brain embeddings to cognitive processes and behavioral measures could enhance our understanding of the cognitive mechanisms underlying language. [32] underscores the need to understand properties encoded in sentence representations, which could inform future research in this area. Addressing these challenges may lead to more cognitively plausible language models and deeper insights into the neural basis of language.
@@ V-A1 Need for Compression: Computational and Storage Challenges of Large Embedding Matrices
The impressive performance of large language models (LLMs) in NLP applications often comes with substantial computational and storage requirements, particularly due to the embedding layers that represent vocabulary items as dense vectors. As vocabularies expand to include hundreds of thousands of tokens and embedding dimensions increase to capture nuanced semantic relationships, the memory footprint can become prohibitive, especially in resource-constrained environments like mobile devices. This scalability challenge is further compounded by the increasing demand for multilingual and multimodal models, which necessitate even larger embedding matrices to represent diverse linguistic and perceptual information [35, 29]. Efficient compression techniques are essential for effectively deploying these powerful models [76]. Additionally, the computational cost associated with large embedding matrices can impede training and inference speed, impacting real-time applications [10, 26].
@@ V-A2 Compression Techniques
Knowledge distillation trains a smaller ”student” model to replicate the behavior of a larger ”teacher” model, effectively transferring knowledge from a complex model to a more efficient one. In cross-lingual embeddings, this technique helps transfer the rich semantic space of a high-resource language model to a smaller model that supports multiple languages, as demonstrated in [52]. This approach is particularly valuable for low-resource languages by leveraging knowledge from data-rich languages [28]. Knowledge distillation can also be applied for vocabulary reduction, enabling a smaller embedding matrix to capture essential semantic information with a condensed vocabulary [77].
@@
Weight tying reduces model size by sharing parameters between different layers, commonly between input and output embedding matrices in language models [46, 78]. This approach not only decreases the memory footprint but can also enhance performance by promoting consistency between input and output representations. [79] further proposes a modification that decouples the hidden state from word prediction, achieving comparable or better results with fewer parameters.
@@
Quantization reduces the precision of embedding vectors (e.g., using 8-bit integers instead of 32-bit floats), significantly lowering storage requirements and accelerating computations, though potentially at the cost of precision loss. Quantization techniques are particularly effective for memory-constrained devices [80].
@@
Pruning techniques reduce model size by removing redundant or less important dimensions or vectors from the embedding matrix. By carefully selecting pruning criteria, this method can maintain performance while making the model more efficient, though excessive pruning can degrade performance [81].
@@
Subspace embedding methods map high-dimensional embeddings into lower-dimensional subspaces. For instance, [82] uses a Cartesian product of subspaces to construct compact embeddings, while [83] applies Tensor-Train Decomposition (TTD) to compress embedding layers, achieving significant size reduction with minimal or no performance loss. Additionally, [84] projects word history vectors onto multiple diverse low-dimensional subspaces.
@@ V-A3 Evaluating Compression Methods: Balancing Compression Rate and Task Performance
Evaluating embedding compression techniques involves balancing the compression rate against performance on downstream tasks. High compression rates improve efficiency but can adversely impact model accuracy. [77] demonstrated significant compression with minimal performance drop, while [82] achieved high compression with only a slight decrease in accuracy. Selecting an appropriate compression method requires careful consideration of the target application and tolerance for performance degradation.
@@ V-B1 The Black Box Nature of Embedding Spaces
Embedding models have revolutionized natural language processing by capturing semantic relationships between words, phrases, and sentences. However, the internal mechanisms of these models remain largely opaque, creating challenges for interpretability. Although embeddings perform well in various tasks, understanding why they work and what specific information they encode remains difficult. This ”black box” nature limits our ability to fully trust and effectively debug these models. Traditional word embeddings, for instance, can struggle with nuanced word-level relationships [55]. Transformer-based models, despite excelling at complex NLP tasks, present additional interpretability challenges due to their depth and complexity, making it difficult to understand the properties encoded in sentence representations and potential risks of information leakage [32]. As models grow more sophisticated, developing techniques to interpret their embeddings becomes increasingly critical.
@@ V-B2 Methods for Understanding Embeddings
Visualizing embedding spaces provides insights into relationships between words, sentences, and concepts. Dimensionality reduction techniques like Principal Component Analysis (PCA), t-distributed Stochastic Neighbor Embedding (t-SNE), and Uniform Manifold Approximation and Projection (UMAP) are often used to project high-dimensional embeddings into two or three dimensions for visualization. For example, [85] visualized word embeddings to understand how Arabic word semantics are represented. Similarly, [86] applied dimensionality reduction to visualize the semantic space of text in the sports domain. However, visualizations come with limitations: projecting high-dimensional data into lower dimensions involves information loss, and the choice of technique can affect the insights gained.
@@
Probing tasks are designed to evaluate specific linguistic properties within embeddings. These tasks involve training a simple classifier (e.g., linear classifier) to predict linguistic attributes (e.g., part-of-speech, tense) from embeddings, with classifier performance indicating the extent to which embeddings capture these attributes. For instance, [87] used probing tasks to compare the linguistic knowledge in BERT and Word2vec embeddings. Additionally, [88] analyzed how word representation varies across layers in neural language models, finding that deeper layers capture complex linguistic features like semantics and morphology. [55] further explored how input and output embeddings in neural models encode different semantic knowledge.
@@
Concept Activation Vectors (CAVs) offer insights into the contribution of specific concepts to a model’s predictions. A CAV represents a direction in the embedding space corresponding to a particular concept. By projecting embeddings onto this direction, we can measure the association of embeddings with that concept. [89] briefly references CAVs but does not provide specifics on their implementation.
@@
Embedding inversion attacks aim to reconstruct the original text from its embedding, highlighting privacy risks and potential information leakage. [90] proposed a generative inversion attack that could reconstruct input sequences from sentence embeddings, exposing the possibility of recovering sensitive information from anonymized representations.
@@ V-B3 Using LLMs to Interpret Embeddings
Large language models (LLMs) present a novel approach to interpreting embeddings by generating explanations and narratives for embedding vectors. As demonstrated in [89], LLMs can translate abstract embeddings into understandable descriptions, enhancing interpretability and expanding applications for embeddings.
@@ V-B4 Applications of Explainable Embeddings
Explainable embeddings are essential for developing more trustworthy AI systems and improving model debugging. By understanding the encoded information and its impact on predictions, we can identify and mitigate biases [91]. Explainable embeddings also aid in debugging by clarifying why a model made specific predictions, allowing identification and correction of errors or biases in the model’s training data or architecture. While interpretability remains a challenge, ongoing research in methods like visualization, probing tasks, CAVs, embedding inversion attacks, and leveraging LLMs for interpretation, are paving the way for more transparent and reliable language models.
@@ V-C1 Challenges of Representing Numbers in Text
Representing numbers poses unique challenges for language models due to their inherent quantitative properties, which differ from typical lexical tokens. Unlike words, numbers have magnitudes and relationships that are essential for understanding quantitative context. For example, the difference between ”1” and ”2” is a quantitative one, not just a lexical difference, and the relationship between ”1” and ”10” is distinct from that between ”9” and ”10,” despite similar lexical distances. Standard word embedding methods often treat numbers as regular tokens, failing to capture these crucial quantitative and relational aspects. This limitation becomes particularly problematic for tasks requiring numerical reasoning or quantitative information extraction, where the model must understand magnitudes and relationships between numbers [92].
@@ V-C2 Embedding Numerical Information
In most language models, numbers are treated as standard tokens and are assigned vector representations similar to words. This approach does not inherently encode the quantitative properties of numbers, often resulting in poor performance on tasks involving numerical magnitudes or relationships [42]. The limitations of this method are most apparent in mathematical reasoning and quantitative information extraction tasks, where an understanding of numbers is crucial.
@@
Exponential embeddings seek to address these limitations by encoding numerical magnitudes in the embedding space, often using a logarithmic scale. This approach allows the embedding distance to approximate the ratio between numbers, capturing some aspects of numerical magnitude. However, while exponential embeddings offer an improvement over default methods, they still struggle to fully represent numerical relationships [37].
@@
Floating-point embeddings take inspiration from the computer’s floating-point representation, encoding numbers as a combination of a mantissa and an exponent. This enables models to capture both the magnitude and precision of numerical values, improving their comprehension of numerical information and enhancing performance on numerical reasoning tasks [93].
@@ V-C3 Specialized Architectures for Numerical Reasoning in LLMs
To bolster numerical reasoning capabilities in large language models (LLMs), researchers are developing specialized architectures that integrate external knowledge sources or modules for handling numerical and mathematical operations. Such architectures provide promising advancements for tasks requiring numerical reasoning, quantitative information extraction, and other forms of numerical understanding [94].
@@ V-C4 Evaluation of Numerical Reasoning Capabilities: Datasets and Metrics
Evaluating the numerical reasoning abilities of language models requires specialized datasets and metrics. Datasets like Numeracy-600K and Wiki-Concert contain textual data with numerical information, serving as benchmarks to assess model performance on numerical reasoning tasks [93]. Key metrics—such as accuracy, precision, recall, and F1-score—help quantify a model’s capability to correctly extract and process numerical information. These datasets and metrics are crucial for tracking advancements in developing language models with strong numerical reasoning capabilities.

### v2

@@ IV-C2 Methods for Modeling Brain Activity
Several neuroimaging techniques allow researchers to study brain activity during language processing. Functional magnetic resonance imaging (fMRI) measures brain activity through blood flow changes, providing insight into the brain regions involved in various language tasks. [shailee_incorporating_2018] used fMRI to analyze the impact of contextual information on brain responses. Electroencephalography (EEG) captures electrical activity from scalp electrodes, offering high temporal resolution for tracking neural response timing to language stimuli. [andrea_family_2023] applied EEG to examine brain responses to familiar entities. Electrocorticography (ECoG), which involves placing electrodes on the brain’s surface, provides even higher spatial resolution than EEG. [ariel_brain_2022] utilized ECoG to study the geometric similarity between brain embeddings and contextual language model embeddings. These techniques capture neural responses to linguistic stimuli, ranging from single words to complex narratives.
@@
Brain embeddings are continuous vector representations derived from neural activity in response to linguistic stimuli. These embeddings encapsulate patterns of brain responses across specific time windows and regions of interest (ROIs). For example, [ariel_brain_2022] extracted brain embeddings by averaging neural activity in the inferior frontal gyrus (IFG) during a listening task. This enables direct comparison with word embeddings derived from language models, facilitating a deeper understanding of how the brain processes language.
@@
Techniques such as geometric alignment and zero-shot mapping are used to compare brain embeddings with language model embeddings. Geometric alignment assesses the structural similarity between embedding spaces, while zero-shot mapping involves predicting brain activity patterns for new words based on their language model representations, or vice versa. [richard_low-dimensional_2021] adapted a transfer learning approach to reveal a low-dimensional structure in language model representations that aligns with brain responses. [ariel_brain_2022] found that brain embeddings and DLM embeddings share similar geometry, enabling zero-shot predictions of brain activity for novel words, suggesting a potential alignment between human and artificial language representation.
@@ IV-C3 Relating Language Model Layers to Temporal Dynamics of Brain Activity
Research has explored the relationship between the layered architecture of deep language models (DLMs) and the temporal dynamics of brain activity. [goldstein_correspondence_2022] found a correspondence between sequential transformations in DLM layers and the temporal sequence of neural activity in brain regions associated with language. This suggests that different DLM layers may reflect different stages of language processing in the brain. However, the study also identifies unique neural processes, such as recurrent information accumulation and responses to unpredictable words, which are not fully captured by current DLMs.
@@
While language models are typically trained on large general corpora, individual experiences shape human language understanding. [andrea_family_2023] demonstrated that language models could represent subject-specific semantic knowledge of familiar people and places when trained on personalized data. This capability opens up possibilities for personalized language models that capture both general and individual semantic knowledge, with applications in personalized recommendations and user-specific content generation.
@@
Despite recent advancements, challenges remain in modeling brain activity with language model embeddings. Understanding the dimensionality of brain embedding spaces and how they relate to language model dimensions is essential. Another key challenge is modeling neural processes for unpredictable words, which current models do not fully capture. Additionally, linking brain embeddings to cognitive processes and behavioral measures could enhance our understanding of the cognitive mechanisms underlying language. [yossi_analysis_2017] underscores the need to understand properties encoded in sentence representations, which could inform future research in this area. Addressing these challenges may lead to more cognitively plausible language models and deeper insights into the neural basis of language.
@@ V-A1 Need for Compression: Computational and Storage Challenges of Large Embedding Matrices
The impressive performance of large language models (LLMs) in NLP applications often comes with substantial computational and storage requirements, particularly due to the embedding layers that represent vocabulary items as dense vectors. As vocabularies expand to include hundreds of thousands of tokens and embedding dimensions increase to capture nuanced semantic relationships, the memory footprint can become prohibitive, especially in resource-constrained environments like mobile devices. This scalability challenge is further compounded by the increasing demand for multilingual and multimodal models, which necessitate even larger embedding matrices to represent diverse linguistic and perceptual information [shao-yen_multimodal_2019, fangxiaoyu_language-agnostic_2020]. Efficient compression techniques are essential for effectively deploying these powerful models [dan_fighting_2018]. Additionally, the computational cost associated with large embedding matrices can impede training and inference speed, impacting real-time applications [a_learning_2013, mourad_word_2022].
@@ V-A2 Compression Techniques
Knowledge distillation trains a smaller ”student” model to replicate the behavior of a larger ”teacher” model, effectively transferring knowledge from a complex model to a more efficient one. In cross-lingual embeddings, this technique helps transfer the rich semantic space of a high-resource language model to a smaller model that supports multiple languages, as demonstrated in [nils_making_2020]. This approach is particularly valuable for low-resource languages by leveraging knowledge from data-rich languages [trevor_cross-lingual_2017]. Knowledge distillation can also be applied for vocabulary reduction, enabling a smaller embedding matrix to capture essential semantic information with a condensed vocabulary [sanqiang_extreme_2019].
@@
Weight tying reduces model size by sharing parameters between different layers, commonly between input and output embedding matrices in language models [hakan_tying_2016, ofir_using_2016]. This approach not only decreases the memory footprint but can also enhance performance by promoting consistency between input and output representations. [kristina_how_2018] further proposes a modification that decouples the hidden state from word prediction, achieving comparable or better results with fewer parameters.
@@
Quantization reduces the precision of embedding vectors (e.g., using 8-bit integers instead of 32-bit floats), significantly lowering storage requirements and accelerating computations, though potentially at the cost of precision loss. Quantization techniques are particularly effective for memory-constrained devices [tao2022compression].
@@
Pruning techniques reduce model size by removing redundant or less important dimensions or vectors from the embedding matrix. By carefully selecting pruning criteria, this method can maintain performance while making the model more efficient, though excessive pruning can degrade performance [wang2019structured].
@@
Subspace embedding methods map high-dimensional embeddings into lower-dimensional subspaces. For instance, [amit_lightweight_2023] uses a Cartesian product of subspaces to construct compact embeddings, while [xu_tensorgpt_2023] applies Tensor-Train Decomposition (TTD) to compress embedding layers, achieving significant size reduction with minimal or no performance loss. Additionally, [kartik_diverse_2014] projects word history vectors onto multiple diverse low-dimensional subspaces.
@@ V-A3 Evaluating Compression Methods: Balancing Compression Rate and Task Performance
Evaluating embedding compression techniques involves balancing the compression rate against performance on downstream tasks. High compression rates improve efficiency but can adversely impact model accuracy. [sanqiang_extreme_2019] demonstrated significant compression with minimal performance drop, while [amit_lightweight_2023] achieved high compression with only a slight decrease in accuracy. Selecting an appropriate compression method requires careful consideration of the target application and tolerance for performance degradation.
@@ V-B1 The Black Box Nature of Embedding Spaces
Embedding models have revolutionized natural language processing by capturing semantic relationships between words, phrases, and sentences. However, the internal mechanisms of these models remain largely opaque, creating challenges for interpretability. Although embeddings perform well in various tasks, understanding why they work and what specific information they encode remains difficult. This ”black box” nature limits our ability to fully trust and effectively debug these models. Traditional word embeddings, for instance, can struggle with nuanced word-level relationships [steven_analysing_2020]. Transformer-based models, despite excelling at complex NLP tasks, present additional interpretability challenges due to their depth and complexity, making it difficult to understand the properties encoded in sentence representations and potential risks of information leakage [yossi_analysis_2017]. As models grow more sophisticated, developing techniques to interpret their embeddings becomes increasingly critical.
@@ V-B2 Methods for Understanding Embeddings
Visualizing embedding spaces provides insights into relationships between words, sentences, and concepts. Dimensionality reduction techniques like Principal Component Analysis (PCA), t-distributed Stochastic Neighbor Embedding (t-SNE), and Uniform Manifold Approximation and Projection (UMAP) are often used to project high-dimensional embeddings into two or three dimensions for visualization. For example, [dima_comparative_2018] visualized word embeddings to understand how Arabic word semantics are represented. Similarly, [rishabh_comparative_2021] applied dimensionality reduction to visualize the semantic space of text in the sports domain. However, visualizations come with limitations: projecting high-dimensional data into lower dimensions involves information loss, and the choice of technique can affect the insights gained.
@@
Probing tasks are designed to evaluate specific linguistic properties within embeddings. These tasks involve training a simple classifier (e.g., linear classifier) to predict linguistic attributes (e.g., part-of-speech, tense) from embeddings, with classifier performance indicating the extent to which embeddings capture these attributes. For instance, [alessio_contextual_2020] used probing tasks to compare the linguistic knowledge in BERT and Word2vec embeddings. Additionally, [matthew_dissecting_2018] analyzed how word representation varies across layers in neural language models, finding that deeper layers capture complex linguistic features like semantics and morphology. [steven_analysing_2020] further explored how input and output embeddings in neural models encode different semantic knowledge.
@@
Concept Activation Vectors (CAVs) offer insights into the contribution of specific concepts to a model’s predictions. A CAV represents a direction in the embedding space corresponding to a particular concept. By projecting embeddings onto this direction, we can measure the association of embeddings with that concept. [tennenholtz_demystifying_2023] briefly references CAVs but does not provide specifics on their implementation.
@@
Embedding inversion attacks aim to reconstruct the original text from its embedding, highlighting privacy risks and potential information leakage. [li_sentence_2023] proposed a generative inversion attack that could reconstruct input sequences from sentence embeddings, exposing the possibility of recovering sensitive information from anonymized representations.
@@ V-B3 Using LLMs to Interpret Embeddings
Large language models (LLMs) present a novel approach to interpreting embeddings by generating explanations and narratives for embedding vectors. As demonstrated in [tennenholtz_demystifying_2023], LLMs can translate abstract embeddings into understandable descriptions, enhancing interpretability and expanding applications for embeddings.
@@ V-B4 Applications of Explainable Embeddings
Explainable embeddings are essential for developing more trustworthy AI systems and improving model debugging. By understanding the encoded information and its impact on predictions, we can identify and mitigate biases [shiva_evaluating_2023]. Explainable embeddings also aid in debugging by clarifying why a model made specific predictions, allowing identification and correction of errors or biases in the model’s training data or architecture. While interpretability remains a challenge, ongoing research in methods like visualization, probing tasks, CAVs, embedding inversion attacks, and leveraging LLMs for interpretation, are paving the way for more transparent and reliable language models.
@@ V-C1 Challenges of Representing Numbers in Text
Representing numbers poses unique challenges for language models due to their inherent quantitative properties, which differ from typical lexical tokens. Unlike words, numbers have magnitudes and relationships that are essential for understanding quantitative context. For example, the difference between ”1” and ”2” is a quantitative one, not just a lexical difference, and the relationship between ”1” and ”10” is distinct from that between ”9” and ”10,” despite similar lexical distances. Standard word embedding methods often treat numbers as regular tokens, failing to capture these crucial quantitative and relational aspects. This limitation becomes particularly problematic for tasks requiring numerical reasoning or quantitative information extraction, where the model must understand magnitudes and relationships between numbers [t_language_2021].
@@ V-C2 Embedding Numerical Information
In most language models, numbers are treated as standard tokens and are assigned vector representations similar to words. This approach does not inherently encode the quantitative properties of numbers, often resulting in poor performance on tasks involving numerical magnitudes or relationships [andrew_modeling_2015]. The limitations of this method are most apparent in mathematical reasoning and quantitative information extraction tasks, where an understanding of numbers is crucial.
@@
Exponential embeddings seek to address these limitations by encoding numerical magnitudes in the embedding space, often using a logarithmic scale. This approach allows the embedding distance to approximate the ratio between numbers, capturing some aspects of numerical magnitude. However, while exponential embeddings offer an improvement over default methods, they still struggle to fully represent numerical relationships [linda_effect_2019].
@@
Floating-point embeddings take inspiration from the computer’s floating-point representation, encoding numbers as a combination of a mantissa and an exponent. This enables models to capture both the magnitude and precision of numerical values, improving their comprehension of numerical information and enhancing performance on numerical reasoning tasks [xiaoxiao_floating-point_2024].
@@ V-C3 Specialized Architectures for Numerical Reasoning in LLMs
To bolster numerical reasoning capabilities in large language models (LLMs), researchers are developing specialized architectures that integrate external knowledge sources or modules for handling numerical and mathematical operations. Such architectures provide promising advancements for tasks requiring numerical reasoning, quantitative information extraction, and other forms of numerical understanding [qiang_bilstm_2021].
@@ V-C4 Evaluation of Numerical Reasoning Capabilities: Datasets and Metrics
Evaluating the numerical reasoning abilities of language models requires specialized datasets and metrics. Datasets like Numeracy-600K and Wiki-Concert contain textual data with numerical information, serving as benchmarks to assess model performance on numerical reasoning tasks [xiaoxiao_floating-point_2024]. Key metrics—such as accuracy, precision, recall, and F1-score—help quantify a model’s capability to correctly extract and process numerical information. These datasets and metrics are crucial for tracking advancements in developing language models with strong numerical reasoning capabilities.